OpenStudy (anonymous):

simplify 16/2+6-4x3 a) 1 b) 30 c) 2 d) 16 solve 12-x=-2 a) 14 b) 10 c) -14 d) -10

OpenStudy (lalaly):

8+6-12 = 14-12 = 2 so (c) x=12+2=14 so (a)